A simple random sample of size n equals 450 individuals who are currently employed is asked if they work at home at least once per week. Of the 450 employed individuals​ surveyed, 28 responded that they did work at home at least once per week. Construct a​ 99% confidence interval for the population proportion of employed individuals who work at home at least once per week.
Answer #1

The 99% Confidence Interval For Proportions

\hat{p} = 28 / 450 = 0.0622, 1 - \hat{p} = 0.9378, n = 450

The Zcritical (2 tail) for \alpha = 0.01, is 2.576

The Confidence Interval is given by \hat{p} \pm ME, where

ME = Zcritical * \sqrt{\frac{\hat{p}(1-\hat{p})}{n} }= 12.576 * \sqrt{\frac{0.0622*0.9378}{450}} = 0.0294

The Lower Limit = 0.0622 - 0.0294 = 0.0328

The Upper Limit = 0.0622 + 0.0294 = 0.0916

The 99% Confidence Interval is (0.0328 , 0.0916)

_________________________________________

If required to 3 decimal place, the CI is (0.033, 0.092)

If required to 2 decimal place, the CI is (0.03, 0.09)